# Enumerating Finite Braid Group Orbits on $SL_{2}(\mathbb{C})$-Character Varieties

### v1

@@ 1 Introduction
Finite irreducible complex reflection groups, which can be viewed as generalizations of Coxeter groups, were classified by Shephard and Todd in 1954 as belonging to an infinite family $G(m,p,n)$ for each rank $n$ or being one of 34 exceptional cases denoted $G_{3}$ to $G_{37}$ [9]. In view of this classification, Lam-Landesman-Litt [17, Cor. 5.1.4] prove that for $T+1\geq 7$, MCG-finite representations do not come from groups of rank five or greater. Their paper leaves open for analysis the infinite family of rank three and four and most of the exceptional groups of rank three and four, as well as the case when $T+1<7$, which we investigate.
Katz [15] originally introduced middle convolution in the language of perverse sheaves and one of our contributions is implementing computer code to explicitly compute the middle convolution given a tuple of matrices, following the algebraic reformulation of Dettweiler-Reiter [6, Section 2]. We also provide code for computing the braid group orbit for MCG-finite 4-tuples and 5-tuples in the $SL_{2}(\mathbb{C})$-character variety of the punctured sphere. This computer code is publicly accessible online at the following link: amalvayalin.github.io/research.html and we hope it encourages further applications of this operation.
@@
With this definition in mind, our problem of studying rank 2 MCG-finite tuples changes to studying nice tuples in the finite complex reflection groups. Using properties of complex reflection groups, as well as prior work on this topic, it remains to study nice $n$ and $n+1$-tuples in the rank $n=3,4$ primitive groups, nice $n+1$-tuples in the imprimitive groups of rank $2<n\leq 5$ and nice $n$-tuples in imprimitive groups of rank $n\geq 3$. More details on why these are the only remaining cases is provided in Remark 3.0.4, Remark 4.0.5 and [17, Cor. 5.1.4].
@@ 1.1 Prior Work
MCG-finite representations correspond to algebraic solutions to nonlinear differential equatons, such as the Painlevé VI equation and more generally the Schlesinger equations [4]. Regarded as a nonlinear analogue of the hypergeometric equation, the Painlevé VI (PVI) equation posed over 100 years ago, has inspired much work on the study of finite orbits. Algebraic solutions to the PVI equation correspond to irreducible MCG-finite representations when $n=2$ and $T+1=4$ and motivates progress towards a nonlinear Schwartz list [2]. Among the many advances made in this area, notable contributions on finite orbits come from Hitchin [14], Dubrovin-Mazzocco [9], Boalch [1], Kitaev [16] and an eventual computer-aided classification by Lisovy-Tykhyy [19].
@@
The connection between finite complex reflection groups and middle convolutions was further explored more recently in the work of Lam-Landesman-Litt [17, Cor. 1.1.7] who proved that all MCG-finite tuples with at least one matrix of infinite order are either pullback representations (classified in [8]) or come from finite complex reflection groups via middle convolution.
Our work investigates the finite complex reflection groups for MCG-finite tuples and examines their middle convolution in relation to the braid group orbit. In the applicable cases, we also explore how the examples we find relate to the classifications provided by Lisovy-Tykhyy [19] and Tykhyy [27]. Note that the analysis in these two papers exclude the tuples where all the matrices share a common eigenvector; these correspond to non-irreducible cases and were classified by Cousin-Moussard [11, Section 3].
@@ 2.1 Middle Convolution
Letting $k$ be a field and $F_{T}$ denote the free group on $T$ generators, middle convolution can equivalently be understood as a functor on the category of finite dimensional $k[F_{T}]$ modules to itself as described by Dettweiler-Reiter [5, Prop. 2.6]. The relation to local systems on the $T+1$-punctured sphere is by taking $k=\mathbb{C}$ and viewing a tuple of $T$ matrices as defining the action of $\mathbb{C}[F_{T}]$ on the module i.e. a representation of $\mathbb{C}[F_{T}]$. In this way irreducible tuples correspond to irreducible representations. While middle convolution preserves irreducibility [5, Cor. 3.6], it typically changes the dimension of the module (rank of the local system on $T+1$-punctured sphere) and does not preserve basic properties e.g. finiteness of the subgroup generated by the matrices (i.e. monodromy group). Important to the study of finite orbits, one desirable property of middle convolution is that it is equivariant with the action of the braid group $B_{T}$ [5, Theorem 5.1]. One way to construct nontrivial finite orbits is to take irreducible tuples generating a finite subgroup, hence finite braid group orbit, and taking the middle convolution to (hopefully) get a tuple generating an infinite group which will necessarily be irreducible and have finite braid group orbit.
@@
Proofs of this statement, as well as further details and properties of the middle convolution can be found in [6, Section 2] and [5, Section 2-5]. ∎
@@
Proved in [5, Theorem 5.5]. ∎
Our code for middle convolution uses the formulation given by Dettweiler-Reiter in [6, Definition 2.1] as opposed to their formulation in [5, Section 2] which only differs by changing the order of blocks in the matrices. The code requires a matrix group $W$ containing the matrices in the tuple $\mathbf{A}$ and the base ring of $W$ should contain the parameter $\lambda$ used for middle convolution. Note that the final function $Compute\_MC$ will automatically check if our intermediate constructions act invariantly on the subspaces $K$ and $L$, as defined in the notation of [6, Definition of 2.1]. The full computer code can be found online at amalvayalin.github.io/research.html.
@@ 2.2 Finite Complex Reflection Groups
As we want irreducible MCG-finite T+1-tuples, we know that our nice $T$-tuple (see Defintion 3.0.1) must at least be of length $n$ to generate an irreducible FCRG, so in each group we only need to search for $T$-tuples with $T>n$. Among the rank three and four primitive irreducible groups, only $G_{31}$ is not well-generated (see Cartan matrices in [18, Appendix Section 3]. An imprimitive group $G(m,p,n)$ is well-generated if $p=m$ or $p=1$.
Associated to each FCRG of rank $n$ is a set of $n$ invariants called the degrees of the reflection group [22, Theorem 5.1], [18, Table D.3]. An important property of these invariants relevant for our purposes is their relation to the eigenvalues of the group elements. Note that we already know the eigenvalues for any element of the group must be roots of unity since the group is finite.
@@ 2.3 Representations and Character Varieties
For proof and further discussion of character varieties of the punctured sphere see [12, Section 4] and [21, Section 5]. ∎
@@ 3.2 Finding Nice Tuples
By Theorem 2.2.2, an eigenvalue of multiplicity $T-2$ for an element in $W$ must be a $d$-th root of unity for $d$ dividing at least $T-2$ degrees of $W$. Any reflection group can be constructed in Magma by providing the generating reflections; the irreducible finite complex reflection groups can be easily constructed in Magma by providing the Shephard-Todd number or the corresponding parameters for $G(m,p,r)$. In most cases Magma generates the group $W$ via the reflection representation over the trace field of $W$ [10, Table 1, Table 2].
@@ 4 Imprimitive Reflection Groups
Imprimitive reflection groups $G(m,p,n)$ are parameterized by $m,p,n\geq 1$ with $p\mid m$. These groups can be viewed as $n\times n$ generalized permutation matrices, which each row and column containing only one non-zero entry which is either 1 or an $\frac{m}{p}$-th root of unity. For $m>1$ and $(m,p,n)\neq(2,2,2)$ these groups are always irreducible, meaning they act irreducibly on $\mathbb{C}^{n}$ (see [3, Section 2]).
An element $A$ of $G(m,p,n)$ can be written as $A=[a_{1},a_{2},\dots,a_{n}\mid\sigma]$ for $\sigma\in S_{n}$ and $p\mid a_{1}+a_{2}\dots+a_{n}$. This element can be identified with the matrix where the $(i,\sigma(i))$-th entry of $A$ is $\zeta_{m}^{a_{i}}$ and all the other entries are 0 [25, Section 2.3]. Note that all these $a_{i}$ should be read modulo $m$ since we are working with powers of an $m$-th root of unity. Additionally, there is a natural surjective homomorphism $\varphi:G(m,p,n)\longrightarrow S_{n}$ given by sending $A\mapsto\sigma$.
@@
    Type 1: These are elements such that $\sigma=(ij)$ is a transposition with $a_{k}\equiv 0\mod m$ for $k\neq i,j$ and $a_{i}+a_{j}\equiv 0\mod m$. All reflections of Type 1 are of order 2 and lie in the subgroup $G(m,m,n)$ of $G(m,p,n)$. Using the notation from [25, Section 2.3] we write a Type 1 reflection as $s(i,j;a_{i})$ with the understanding that this is the same as $s(j,i;-a_{i})$.
@@
Stated in [25, Lemma 3.2] with proofs in [23, Subsection 2.10 and Lemma 2.1] and [24, Lemma 2.2]. ∎
@@
This is true by the definition of reflections and the definition of the homomorphism. For a more thorough examination of related properties, we refer to [3, Proposition 2.2]. ∎
Given any set of reflections generating an irreducible imprimitive group the permutations associated to the reflections in the set must generate a transitive subgroup of $S_{n}$, or else the group cannot be irreducible. Note that in the case where the associated permutations are transpositions (i.e. the reflections are Type 1), the only transitive subgroup of $S_{n}$ generated by transpositions is $S_{n}$ itself [18, Lemma 2.13].
@@
The first is proven in [23, Thm. 2.19] and the second is proven in [28, Theorem 3.10]. Note that in both these papers they refer to cycles as circles. The third statement is proven in [28, Theorem 3.9], where they define rooted trees to capture the presence of Type 2 reflections. We refer to these paper for more details on the graph theoretic interpretation of generating sets and their connection to reflection subgroups of the imprimitive reflection groups. ∎
@@
Among the imprimitive groups, we have to check for $n$-tuples and $n+1$-tuples in $G(m,p,n)$. By Theorem 4.0.1, the only possible groups that can contain nice $n$-tuples are the well-generated groups $G(m,1,n)$ and $G(m,m,n)$. We prove there are no nice $n$-tuples in either of these groups when $n\geq 5$ in Section 4.5. By results in [17, Lemma 5.3.2], we know that $n+1$-tuples can only exist in $G(m,p,n)$ for $n<5$. This work will fully analyze the existence of nice $n+1$-tuples for $52<n<$ 5.
@@ 5.1 Nice 3-Tuples in Rank 3 Groups
Nice tuples in this case will be tuples of 3 reflections generating the whole group such that their inverse product has some nontrivial eigenvalue of multiplicity 1. After middle convolution, our the induced tuple will correspond to $SL_{2}(\mathbb{C})$ local systems on the 4-punctured sphere and we can compare our braid group orbits to those in the literature. We compute the parameters $(\omega_{X},\omega_{Y},\omega_{Z},\omega_{4})$ in the notation of [19] and found that the $\omega$ parameters for each of our orbits match (up to permutation) with one listed in the work of Lisovy-Tykhyy. Since Tykhyy’s classification [27] further refines the original classification by Lisovy-Tykhy [19], each tuple of $\omega$ parameters gives multiple orbits that Tykhyy identifies uniquely in terms of residues of the trace coordinates $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})$ and $(\sigma_{12},\sigma_{23},\sigma_{13},\sigma_{24})$ for some representative. In the nice 3-tuple case, using the $\omega$ parameters, we are able to easily identify how our orbit size relates the orbit size listed by Tykhyy for each $\mathcal{M}^{(4)}$ (in Tykhyy’s notation) and narrow down the possibilities for which of Tykhyy’s orbits our tuple corresponds to. In Tykhyy’s classification, tuples are considered equivalent up to: multiplying any two matrices by -1, complex conjugation of all elements of all matrices, cyclically permuting the matrices in the tuple, and up to taking inverses and reversing the order of the tuple [27]. We perform all these operations and compute the residue signatures for the matrices in each orbit to find one that matches the example provided in Tykhyy’s work. In most cases we found exact matches, and in other cases we provide the closest match, marking the differing coordinate with an asterisk (see Section 5.1.3 and 5.1.4).
@@ 5.1.1 $G_{23}$
The group $G_{23}$ of order 120 is also denoted $W(H_{3})$ with degrees $\deg(G_{23})=\{2,6,10\}$ [18, Table D.3]. All reflections in this group are conjugate with nontrivial eigenvalue -1. For nice 3-tuples in $G_{23}$, the computation returned three possible sets of eigenvalues for the inverse product.
@@ 5.1.2 $G_{24}$
The group $G_{24}$ of order 336 is also denoted $W(J_{3}(4))$ with degrees $\deg(G_{24})=\{4,6,14\}$ [18, Table D. 3]. All reflections in this group are conjugate with nontrivial eigenvalue -1. Looking at the degrees, we first expand the base field to be the cyclotomic field of degree 84. There is one inverse pair of nice 3-tuples in this group, so we only provide the details for one in Table 3.

### v2

@@ 1 Introduction
Finite irreducible complex reflection groups, which can be viewed as generalizations of Coxeter groups, were classified by Shephard and Todd in 1954 as belonging to an infinite family $G(m,p,n)$ for each rank $n$ or being one of 34 exceptional cases denoted $G_{3}$ to $G_{37}$ [9]. In view of this classification, Lam-Landesman-Litt [20, Cor. 5.1.4] prove that for $T+1\geq 7$, MCG-finite representations do not come from groups of rank five or greater. Their paper leaves open for analysis the infinite family of rank three and four and most of the exceptional groups of rank three and four, as well as the case when $T+1<7$, which we investigate.
Katz [18] originally introduced middle convolution in the language of perverse sheaves and one of our contributions is implementing computer code to explicitly compute the middle convolution given a tuple of matrices, following the algebraic reformulation of Dettweiler-Reiter [9, Section 2]. We also provide code for computing the braid group orbit for MCG-finite 4-tuples and 5-tuples in the $SL_{2}(\mathbb{C})$-character variety of the punctured sphere. This computer code is publicly accessible online at the following link: amalvayalin.github.io/research.html and we hope it encourages further applications of this operation.
@@
With this definition in mind, our problem of studying rank 2 MCG-finite tuples changes to studying nice tuples in the finite complex reflection groups. Using properties of complex reflection groups, as well as prior work on this topic, it remains to study nice $n$ and $n+1$-tuples in the rank $n=3,4$ primitive groups, nice $n+1$-tuples in the imprimitive groups of rank $2<n\leq 5$ and nice $n$-tuples in imprimitive groups of rank $n\geq 3$. More details on why these are the only remaining cases is provided in Remark 3.0.4, Remark 4.0.5 and [20, Cor. 5.1.4].
@@ 1.1 Prior Work
MCG-finite representations correspond to algebraic solutions to nonlinear differential equatons, such as the Painlevé VI equation and more generally the Schlesinger equations [7]. Regarded as a nonlinear analogue of the hypergeometric equation, the Painlevé VI (PVI) equation posed over 100 years ago, has inspired much work on the study of finite orbits. Algebraic solutions to the PVI equation correspond to irreducible MCG-finite representations when $n=2$ and $T+1=4$ and motivates progress towards a nonlinear Schwartz list [3]. Among the many advances made in this area, notable contributions on finite orbits come from Hitchin [17], Dubrovin-Mazzocco [12], Boalch [2], Kitaev [19] and an eventual computer-aided classification by Lisovy-Tykhyy [22]. The history and further details regarding the 45 equivalence classes of exceptional algebraic solutions of PVI can be found in [1]. Research on finite orbits of the character variety is still ongoing, see [25], [5], and more recently [4].
@@
The connection between finite complex reflection groups and middle convolutions was further explored more recently in the work of Lam-Landesman-Litt [20, Cor. 1.1.7] who proved that all MCG-finite tuples with at least one matrix of infinite order are either pullback representations (classified in [11]) or come from finite complex reflection groups via middle convolution.
Our work investigates the finite complex reflection groups for MCG-finite tuples and examines their middle convolution in relation to the braid group orbit. In the applicable cases, we also explore how the examples we find relate to the classifications provided by Lisovy-Tykhyy [22] and Tykhyy [31]. Note that the analysis in these two papers exclude the tuples where all the matrices share a common eigenvector; these correspond to non-irreducible cases and were classified by Cousin-Moussard [14, Section 3].
@@ 2.1 Middle Convolution
Letting $k$ be a field and $F_{T}$ denote the free group on $T$ generators, middle convolution can equivalently be understood as a functor on the category of finite dimensional $k[F_{T}]$ modules to itself as described by Dettweiler-Reiter [8, Prop. 2.6]. The relation to local systems on the $T+1$-punctured sphere is by taking $k=\mathbb{C}$ and viewing a tuple of $T$ matrices as defining the action of $\mathbb{C}[F_{T}]$ on the module i.e. a representation of $\mathbb{C}[F_{T}]$. In this way irreducible tuples correspond to irreducible representations. While middle convolution preserves irreducibility [8, Cor. 3.6], it typically changes the dimension of the module (rank of the local system on $T+1$-punctured sphere) and does not preserve basic properties e.g. finiteness of the subgroup generated by the matrices (i.e. monodromy group). Important to the study of finite orbits, one desirable property of middle convolution is that it is equivariant with the action of the braid group $B_{T}$ [8, Theorem 5.1]. One way to construct nontrivial finite orbits is to take irreducible tuples generating a finite subgroup, hence finite braid group orbit, and taking the middle convolution to (hopefully) get a tuple generating an infinite group which will necessarily be irreducible and have finite braid group orbit.
@@
Proofs of this statement, as well as further details and properties of the middle convolution can be found in [9, Section 2] and [8, Section 2-5]. ∎
@@
Proved in [8, Theorem 5.5]. ∎
Our code for middle convolution uses the formulation given by Dettweiler-Reiter in [9, Definition 2.1] as opposed to their formulation in [8, Section 2] which only differs by changing the order of blocks in the matrices. The code requires a matrix group $W$ containing the matrices in the tuple $\mathbf{A}$ and the base ring of $W$ should contain the parameter $\lambda$ used for middle convolution. Note that the final function $Compute\_MC$ will automatically check if our intermediate constructions act invariantly on the subspaces $K$ and $L$, as defined in the notation of [9, Definition of 2.1]. The full computer code can be found online at amalvayalin.github.io/research.html.
@@ 2.2 Finite Complex Reflection Groups
As we want irreducible MCG-finite T+1-tuples, we know that our nice $T$-tuple (see Defintion 3.0.1) must at least be of length $n$ to generate an irreducible FCRG, so in each group we only need to search for $T$-tuples with $T>n$. Among the rank three and four primitive irreducible groups, only $G_{31}$ is not well-generated (see Cartan matrices in [21, Appendix Section 3]. An imprimitive group $G(m,p,n)$ is well-generated if $p=m$ or $p=1$.
Associated to each FCRG of rank $n$ is a set of $n$ invariants called the degrees of the reflection group [26, Theorem 5.1], [21, Table D.3]. An important property of these invariants relevant for our purposes is their relation to the eigenvalues of the group elements. Note that we already know the eigenvalues for any element of the group must be roots of unity since the group is finite.
@@ 2.3 Representations and Character Varieties
For proof and further discussion of character varieties of the punctured sphere see [15, Section 4] and [24, Section 5]. ∎
@@ 3.2 Finding Nice Tuples
By Theorem 2.2.2, an eigenvalue of multiplicity $T-2$ for an element in $W$ must be a $d$-th root of unity for $d$ dividing at least $T-2$ degrees of $W$. Any reflection group can be constructed in Magma by providing the generating reflections; the irreducible finite complex reflection groups can be easily constructed in Magma by providing the Shephard-Todd number or the corresponding parameters for $G(m,p,r)$. In most cases Magma generates the group $W$ via the reflection representation over the trace field of $W$ [13, Table 1, Table 2].
@@ 4 Imprimitive Reflection Groups
Imprimitive reflection groups $G(m,p,n)$ are parameterized by $m,p,n\geq 1$ with $p\mid m$. These groups can be viewed as $n\times n$ generalized permutation matrices, which each row and column containing only one non-zero entry which is either 1 or an $\frac{m}{p}$-th root of unity. For $m>1$ and $(m,p,n)\neq(2,2,2)$ these groups are always irreducible, meaning they act irreducibly on $\mathbb{C}^{n}$ (see [6, Section 2]).
An element $A$ of $G(m,p,n)$ can be written as $A=[a_{1},a_{2},\dots,a_{n}\mid\sigma]$ for $\sigma\in S_{n}$ and $p\mid a_{1}+a_{2}\dots+a_{n}$. This element can be identified with the matrix where the $(i,\sigma(i))$-th entry of $A$ is $\zeta_{m}^{a_{i}}$ and all the other entries are 0 [29, Section 2.3]. Note that all these $a_{i}$ should be read modulo $m$ since we are working with powers of an $m$-th root of unity. Additionally, there is a natural surjective homomorphism $\varphi:G(m,p,n)\longrightarrow S_{n}$ given by sending $A\mapsto\sigma$.
@@
    Type 1: These are elements such that $\sigma=(ij)$ is a transposition with $a_{k}\equiv 0\mod m$ for $k\neq i,j$ and $a_{i}+a_{j}\equiv 0\mod m$. All reflections of Type 1 are of order 2 and lie in the subgroup $G(m,m,n)$ of $G(m,p,n)$. Using the notation from [29, Section 2.3] we write a Type 1 reflection as $s(i,j;a_{i})$ with the understanding that this is the same as $s(j,i;-a_{i})$.
@@
Stated in [29, Lemma 3.2] with proofs in [27, Subsection 2.10 and Lemma 2.1] and [28, Lemma 2.2]. ∎
@@
This is true by the definition of reflections and the definition of the homomorphism. For a more thorough examination of related properties, we refer to [6, Proposition 2.2]. ∎
Given any set of reflections generating an irreducible imprimitive group the permutations associated to the reflections in the set must generate a transitive subgroup of $S_{n}$, or else the group cannot be irreducible. Note that in the case where the associated permutations are transpositions (i.e. the reflections are Type 1), the only transitive subgroup of $S_{n}$ generated by transpositions is $S_{n}$ itself [21, Lemma 2.13].
@@
The first is proven in [27, Thm. 2.19] and the second is proven in [32, Theorem 3.10]. Note that in both these papers they refer to cycles as circles. The third statement is proven in [32, Theorem 3.9], where they define rooted trees to capture the presence of Type 2 reflections. We refer to these paper for more details on the graph theoretic interpretation of generating sets and their connection to reflection subgroups of the imprimitive reflection groups. ∎
@@
Among the imprimitive groups, we have to check for $n$-tuples and $n+1$-tuples in $G(m,p,n)$. By Theorem 4.0.1, the only possible groups that can contain nice $n$-tuples are the well-generated groups $G(m,1,n)$ and $G(m,m,n)$. We prove there are no nice $n$-tuples in either of these groups when $n\geq 5$ in Section 4.5. By results in [20, Lemma 5.3.2], we know that $n+1$-tuples can only exist in $G(m,p,n)$ for $n<5$. This work will fully analyze the existence of nice $n+1$-tuples for $52<n<$ 5.
@@ 5.1 Nice 3-Tuples in Rank 3 Groups
Nice tuples in this case will be tuples of 3 reflections generating the whole group such that their inverse product has some nontrivial eigenvalue of multiplicity 1. After middle convolution, our the induced tuple will correspond to $SL_{2}(\mathbb{C})$ local systems on the 4-punctured sphere and we can compare our braid group orbits to those in the literature. We compute the parameters $(\omega_{X},\omega_{Y},\omega_{Z},\omega_{4})$ in the notation of [22] and found that the $\omega$ parameters for each of our orbits match (up to permutation) with one listed in the work of Lisovy-Tykhyy. Since Tykhyy’s classification [31] further refines the original classification by Lisovy-Tykhyy [22], each tuple of $\omega$ parameters gives multiple orbits that Tykhyy identifies uniquely in terms of residues of the trace coordinates $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})$ and $(\sigma_{12},\sigma_{23},\sigma_{13},\sigma_{24})$ for some representative. In the nice 3-tuple case, using the $\omega$ parameters, we are able to easily identify how our orbit size relates the orbit size listed by Tykhyy for each $\mathcal{M}^{(4)}$ (in Tykhyy’s notation) and narrow down the possibilities for which of Tykhyy’s orbits our tuple corresponds to. In Tykhyy’s classification, tuples are considered equivalent up to: multiplying any two matrices by -1, complex conjugation of all elements of all matrices, cyclically permuting the matrices in the tuple, and up to taking inverses and reversing the order of the tuple [31]. We perform all these operations and compute the residue signatures for the matrices in each orbit to find one that matches the example provided in Tykhyy’s work. In most cases we found exact matches, and in other cases we provide the closest match, marking the differing coordinate with an asterisk (see Section 5.1.3 and 5.1.4).
@@ 5.1.1 $G_{23}$
The group $G_{23}$ of order 120 is also denoted $W(H_{3})$ with degrees $\deg(G_{23})=\{2,6,10\}$ [21, Table D.3]. All reflections in this group are conjugate with nontrivial eigenvalue -1. For nice 3-tuples in $G_{23}$, the computation returned three possible sets of eigenvalues for the inverse product.
@@ 5.1.2 $G_{24}$
The group $G_{24}$ of order 336 is also denoted $W(J_{3}(4))$ with degrees $\deg(G_{24})=\{4,6,14\}$ [21, Table D. 3]. All reflections in this group are conjugate with nontrivial eigenvalue -1. Looking at the degrees, we first expand the base field to be the cyclotomic field of degree 84. There is one inverse pair of nice 3-tuples in this group, so we only provide the details for one in Table 3.